OpenStudy (anonymous):

5x + 3y = -2. Find the rate of change.

OpenStudy (anonymous):

Rate of change= slope. Get it into point slope form. y=mx+b, where m is the slope.

OpenStudy (anonymous):

change into linear form of y = mx + b, and you'll get y = -5/3 - 2/3 so the m is the slope or the rate of change is -5/3

OpenStudy (anonymous):

missed the x it's supposed to be y = -5/3x - 2/3

OpenStudy (anonymous):

The coefficient on x is the slope.

OpenStudy (anonymous):

Rate Of Change : 5x+3y=-2 subtract 5x from both sides 5x-5x + 3y = -2 -5x 3y = -5x - 2 Divide both sides by 3 3y/3 = (-5x - 2)/3 y = -5/3x - 2/3

OpenStudy (anonymous):

mk. I'm still a little confused though. Why is the answer y = -5/3x - 2/3?

OpenStudy (anonymous):

It's the point slope form of the equation y = m (slope or rate of change) x + b (y-intercept)

OpenStudy (anonymous):

alright. thanks.

OpenStudy (anonymous):

You have to solve the equation for Y and you'll get the point slope form...